OpenStudy (anonymous):

Prove: cos(theta) = 1/2[e^i(theta) + e^(-i(theta)) I know we use Eulers relation: e^i(theta) = cos(theta) + jsin(theta) so far this is what I get: cos(theta) = e^i(theta) - isin(theta)

OpenStudy (caozeyuan):

so the Q is \[\cos \theta =\frac{ e^{i \theta}+e^{-i \theta}}{ 2 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont know how to write it like that...

OpenStudy (caozeyuan):

and you have to know that cos is an even function and sin is an odd function

OpenStudy (caozeyuan):

which means \[\cos(-\theta) = \cos \theta, \sin(-\theta)=-\sin(\theta)\]

OpenStudy (caozeyuan):

so \[e ^{-i \theta}=\cos(\theta)- isin(\theta)\]

OpenStudy (caozeyuan):

while\[e^{i \theta}=\cos(\theta)+isin(\theta)\]

OpenStudy (anonymous):

oh I think I know where your going with this.

OpenStudy (caozeyuan):

if you add this two together you will get 2 cos theta, and a half of this is the LHS

OpenStudy (anonymous):

yes ok thats what I thought Im just doing the math now. Thanks!

OpenStudy (caozeyuan):

BTW, can you look at my problem? I know the method but get stack with the calculation

OpenStudy (anonymous):

is it posted? sure